### v1

# Verbalized Algorithms

Abstract
Instead of querying LLMs in a one-shot manner and hoping to get the right answer for a reasoning task, we propose a paradigm we call verbalized algorithms (VAs), which leverage classical algorithms with established theoretical understanding. VAs decompose a task into simple elementary operations on natural language strings that they should be able to answer reliably, and limit the scope of LLMs to only those simple tasks. For example, for sorting a series of natural language strings, verbalized sorting uses an LLM as a binary comparison oracle in a known and well-analyzed sorting algorithm (e.g., bitonic sorting network). We demonstrate the effectiveness of this approach on sorting and clustering tasks.
1 Introduction
Imagine receiving thousands of customer complaints and having to quickly sort them by how urgently they need support. Feeding such a large set of input strings blindly into an LLM is ineffective and impractical, both due to its limited context length and its limited ability to perform sorting tasks. LLMs provide no guarantees on the correctness of their outputs for computational reasoning tasks, such as sorting, beyond anecdotal, empirical observations that they sometimes do.
In contrast, traditional disciplines of computer science often approach such a task using task decomposition. For example, a merge sort decomposes the entire sorting task into several sorting tasks over individual sublists and combines the results, where, at the limit, there is an elementary binary comparison operator that compares individual pairs of objects in the list. These traditional algorithms, given an accurate oracle, guarantee the correctness of their outputs and the average/worst-case runtime.
In this work, we propose an abstract framework called “verbalized algorithms”. Given an informal representation of a formal task, a verbalized algorithm operates by replacing some basic atomic operations in a classical algorithm with LLM-driven text-based ones while retaining the high-level control flow of the initial algorithm, thereby largely retaining the theoretical guarantees of the original algorithm.
This has advantages over another common approach to similar tasks that we call the “formalization-based approach”. Leveraging the machine translation ability of LLMs, it converts an informal task representation into a formal, symbolic representation, solves it with a specialized combinatorial solver, and then converts the solution back to the informal, natural language representation for the user. Existing work based on this approach includes LLM-based generation of PDDL (McDermott, 2000) for classical planning (Xie, 2020; Guan et al., 2023; Li et al., 2024; Fine-Morris et al., 2024; Oswald et al., 2024), Prover9 input language for theorem proving (Olausson et al., 2023), or SMT solver input for optimization (Hao et al., 2024). The issue with this approach is twofold: First, it requires grammar-following capability for a complex target formal language. Second, while the downstream solver requires a semantically and logically accurate input, the vagueness and informality in the natural language input often cause LLMs to generate inaccurate and inconsistent translations.
2 Verbalized Algorithms
While whether LLMs can reason or not is up for both philosophical and scientific debates, it is more widely accepted that they can solve some much simpler tasks with high fidelity, such as text-based classification and question-answering. Assuming those capabilities, we propose verbalized algorithms (VAs), an LLM-based reasoning strategy that limits the scope of the LLM’s usage to simple tasks.
Definition 1 (Verbalized Algorithms).
Let $T[\tau]$ be a computational task (e.g., sorting) over objects of type $\tau$ and let $A[\tau]$ be an algorithm that solves $T[\tau]$ with an $n$-arg boolean oracle $f[\tau]:\tau,\ldots,\tau\rightarrow{\left\{\top,\bot\right\}}$. A Verbalized Algorithm instantiates $A[\tau]$ with $\tau=\texttt{str}$ by implementing $f[\texttt{str}]$ as a query to an LLM whose output is constrained to $\left\{\text{``yes''},\text{``no''}\right\}$} which maps to $\left\{\top,\bot\right\}$}.
VAs have three categories: Naive VAs, which treat the LLM’s answer as ground truth, Robust VAs, which treat it as a noisy oracle with a certain error rate and quantify the rate of error in the final output, and Probabilistic VAs, which use the probability/logits of LLM outputs to explicitly quantify the uncertainty of the result and return a distribution of outputs. In this workshop paper, we focus on the naive and the robust variants due to space, leaving the probabilistic variants for future work.
2.1 Verbalized Sorting
Verbalized Sorting sorts a list of natural language strings $X$ using LLM’s question answering capability as an atomic comparison oracle. Each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query such as “Is X better than Y? X: $x_{1}$ Y: $x_{2}$”, where the next output token is restricted to “yes” and “no”. The model output is mapped to $\top/\bot$ if the sampled output is “yes” / “no”, respectively. We evaluated the accuracy of several non-VA and VA approaches. Method, dataset, and prompt details are available in the appendix Sec. A. All results are averages of 5 random seeds.
Constraint Decoding Baseline encodes the list $X$ in a markdown format and asks the model to return a sorted list using structured-decoding (Willard and Louf, 2023) to constrain the output to be a list whose elements are guaranteed to be one of the input strings. Due to the lack of a computationally efficient way to constrain the list to (1) be free of duplicates and (2) represent the same set of strings as the input, we (1) remove the duplicates and (2) append missing elements to the end of the list.
Scoring-based approaches ask the model to assign a score between 1.0 and 5.0 (down to one decimal) to each sentence using regex-constrained decoding, then sort the sentences using the scores. It is a form of formalization-based approach (the list of scores represents the formal problem). We tested two variants: I.i.d Scoring scores $x_{i}\in X$ independently, and Autoregressive Scoring scores all strings in a single query, allowing the model to see the scores of other elements.
VA Powersort overrides Python’s __gt__ operator with an LLM call and applies the built-in sorted function, which uses Powersort (Munro and Wild, 2018) that runs $O(n\log n)$ comparisons worst-case, the best runtime complexity achievable by comparison sort. VA Bitonic sorting network uses the Bitonic sorting network (Batcher, 1968) from parallel sorting literature to improve the throughput. It performs $O(n(\log n)^{2})$ comparisons, but runs in $O((\log n)^{2})$ time due to parallelism. VA Robust Powersort/Bitonic Sorting Network makes individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting. For an odd $K$, if the oracle is correct with probability $p$, the voting result is correct with at least probability $1-e^{-K(2p-1)^{2}/2}$ (Emamjomeh-Zadeh and Kempe, 2018) due to Hoeffding bound (Hoeffding, 1963).
We evaluated the approaches above on Amazon reviews benchmark (Hou et al., 2024). We selected 30 products from the “Video Games” category, which, for each integer rating between 1 and 5, have at least 5 non-empty reviews that are voted as “helpful” by other customers. The task is to sort a shuffled list of 25 selected reviews (5 for each rating) in increasing order of positive sentiment. To measure the accuracy, we compute the Kendall-Tau score (Kendall, 1938) between the output list $Y$ and the ground-truth output list $Y^{*}$, which is sorted by the true rating, averaged over the dataset. We used Qwen3-1.7B and Qwen3-32B-AWQ models (Team, 2025) with thinking enabled.
First, Table 1 (1) shows how often the constraint-decoding baseline violates the sorting invariance, i.e., how many elements are duplicated or missing in the output on average. While these violations become rarer as the model gets larger, they do not disappear. This highlights the lack of theoretical guarantees in LLM-based reasoning, which does not occur in VAs and formalization-based approaches.
Next, Table 1 (2) shows the average Kendall-Tau scores. Constraint-decoding baseline performed the worst overall, and while the scoring approaches can work with a large 32B model, their performance dwindles when the model is smaller. Robust VA with 32B achieved the best results. The strength of VA is most pronounced when using the small 1.7B model, where other approaches failed. Notably, VAs with a 1.7B model outperformed the baseline with a 32B model, clearly showing the advantage of using the sorting algorithms developed in the literature.
Finally, Table 2 compares the runtime of each approach. VA Bitonic Sorting Network improves over the VA Powersort, and both VAs can be faster than the constrained-decoding baseline, even with robustification (x3 parallel queries). This shows that VAs can directly benefit from the theoretical results in the parallel sorting literature, in contrast to the recent LLM reasoning literature that tends to detach from existing discrete algorithm studies. Note that the parallel speedup in bitonic sort does not reach the theoretical limit (x11.4; 171 comparators and 15 steps for 25 inputs) because the maximum batch size is limited by the GPU; multi-GPU experiments with data-parallelism would improve this.
Table 1: (1) Average number of sorting invariance violations by the constraint decoding baseline. “Duplicates” indicates the number of duplicates found in the output, and “missing” indicates the number of input elements that are missing in the output, for 25 input strings. (2) Average Kendall-Tau scores. Top-3 are highlighted in bold. The score of 1, 0, and -1 indicates perfect rank correlations, no correlations, and perfect anti-correlations, respectively.
(1) Invariance (2) Accuracy (Kendall-Tau scores) Models Violations Constraint Scoring Naive VA Robust VA ($K=3$) (Qwen3) duplicates missing Decoding i.i.d. AR Bitonic Powersort Bitonic Powersort 1.7B 1.39 9.63 0.00 0.18 0.05 0.30 0.33 0.33 0.44 32B-AWQ 0.49 2.75 0.30 0.57 0.39 0.39 0.56 0.39 0.58
Table 2: Average total elapsed time in seconds for processing all sorting queries. Bitonic sorting network also shows the relative speedup over the sequential sort (Powersort).
Models Scoring Naive VA Robust VA ($K=3$) (Qwen3) Baseline i.i.d. AR Bitonic Powersort Bitonic Powersort 1.7B 7468 630 205 2338 (x2.12) 4975 5060 (x1.45) 7362 32B-AWQ 7280 1430 626 8556 (x2.18) 18679 20350 (x1.30) 26471
2.2 Verbalized Clustering
Verbalized Clustering attempts to cluster a set of strings $S$ into a family of subsets $\mathcal{S}$ based on their similarity. The problem with text-based clustering is that the distance function for natural language strings is difficult to define. All approaches to this task are embedding methods. Method, dataset, and prompt details are available in the appendix Sec. B. All results are averages of 5 random seeds.
Jaccard Baseline creates a distance matrix using Jaccard distance of token sets of strings, then applies sklearn’s MDS (Kruskal, 1964) to create an embedding. Sentence Transformer (ST) Baseline uses Qwen3-8B-Embedding directly, which was selected as the top performer in the MTEB leaderboard (Muennighoff et al., 2022). VA t-STE: This is a VA based on Student-t Stochastic Triplet Embedding (Van Der Maaten and Weinberger, 2012, t-STE), which makes triplet comparison queries to LLM. Triplet comparison takes three elements $x$, $y$ and $z$ in a metric space $\|\cdot\|$ and query if $\|x-y\|<\|x-z\|$. We ask the LLM a question such as “Is X closer to Y than is to Z? X: $x$, Y: $y$, Z: zz”. For each $x\in S$, we randomly sample $k$ queries from $S^{3}$ and their results from an LLM, then generate $D$-dim embeddings by minimizing the number of violations to the results. We show the results with hyperparameters $k\in{\left\{100,200\right\}},D=8$.
We evaluated the methods across four datasets. Medarxiv (Muennighoff et al., 2022) contains the titles and the categories of medical papers on MedArxiv. The category forms a cluster. IT Field Defects consists of $78$ technical IT field defect reports submitted to our internal defect tracking system (unpublished), where some reports are manually marked as duplicates of another, forming a cluster. xkcdcolors-h/l are two datasets of 100 colors sampled from (Munroe, 2010). Each entry has a non-standard color name (e.g., baby blue, vomit) and a corresponding RGB value. The difference between h and l variants is that we defined 6 ground truth clusters with sklearn.AgglomerativeClustering using a lexicographic distance that prioritizes hue and lightness in the HSL color model, respectively. Assuming that colors tend to be more strongly associated with hues, xkcdcolors-l represents an out-of-distribution (OOD) task.
We evaluate the embeddings with a score defined in Eq. 1, which is the average rank of the members of the cluster each string belongs to, when sorted by the Euclidean distance. The score is normalized to $[0,1]$, where 1 indicates that cluster members are embedded closer. Given the input $S$ and a string $s\in S$, let $C(s)$ be the ground-truth cluster such that $s\in C(s)$. Let $i=\textsc{rank}(s,s^{\prime})\in{\left\{0,1,\ldots|S|-1\right\}}$ be an integer such that, when sorted by the pairwise Euclidean distance $d(e(s),e(s^{\prime\prime}))$ between $s$ and a string $s^{\prime\prime}$ across $s^{\prime\prime}\in S$ using embedding $e$, then $d(e(s),e(s^{\prime}))$ ranks $i$-th. The score is defined as:
score $=\textstyle\quad\frac{1}{|S|}\sum_{s\in S}\frac{1}{|C(s)|-1}\sum_{s^{\prime}\in C(s)\setminus{\left\{s\right\}}}1-\frac{\textsc{rank}(s,s^{\prime})}{|S|}\in[0,1]$ (1)
Table 3 shows the results. Jaccard fared surprisingly well in IT Defects, where the input often contains long output logs whose tokens carry significant information, but it underperformed in tasks that require semantic comprehension. Sentence Transformer achieves strong performance in the standard clustering tasks such as Medarxiv and IT Defects. However, it fell short in xkcdcolors-h/l that requires information about the distance criteria. The score 0.53 in the xkcdcolors-l is close to 0.5, i.e., its distance ranking is only marginally better than a dice roll. VA t-STE outperformed traditional baselines with $k=200$. We highlight its strong generalization in OOD tasks where other approaches failed, due to its strong algorithmic bias combined with flexible LLM queries.
Table 3: Clustering scores across methods and benchmarks. The top 2 results are highlighted in bold.
method model $D=$ $k=$ Medarxiv IT Defects xkcdcolors-h xkcdcolors-l Jaccard 8 0.556 0.767 0.601 0.472 ST Qwen3-Emb…d-8B 4096 0.827 0.936 0.718 0.531 triplet Qwen3-4B 8 100 0.831 0.831 0.823 0.704 triplet Qwen3-4B 8 200 0.858 0.895 0.823 0.691 triplet Qwen3-32B-AWQ 8 100 0.842 0.908 0.851 0.748 triplet Qwen3-32B-AWQ 8 200 0.858 0.939 0.850 0.742
3 Conclusion, Related Work, Discussion and Future Work
We presented Verbalized Algorithms (VAs), a disciplined paradigm for performing algorithmic reasoning with LLMs. It benefits from a strong prior in the golden classics of discrete algorithms published half a century ago, significantly improving the reasoning accuracy even with a small model. It also leverages their established theoretical results to improve the runtime of such reasoning.
Despite significant recent findings in reasoning models, which combine RL-based self-training with “think tags” (Guo et al., 2025), they rarely explicitly perform systematic backtracking and search. Whether such systems can reason or not has been hotly debated in the context of LLM-based planning (Valmeekam et al., 2023). Previous work on solving planning tasks using LLM (Liu et al., 2023; Hirsch et al., 2024) demonstrates varying degrees of success and limitations (e.g., failure in mystery due to sensitivity to the vocabulary). Thought-of-Search (Katz et al., 2024) simulates a systematic search algorithm, which could be seen as an instance of VAs.
While this paper explored only a limited number of applications due to space, we are already pursuing several exciting verbalizations, including submodular maximization and pareto-front computation for diverse / non-dominated document retrieval, DPLL for combinatorial reasoning, and median/top-$k$ selection, etc. The potential of this paradigm is large, and we invite various ideas for unique verbalizations and new applications. For example, can Floyd’s tortoise-and-hare cycle detection algorithm for linked lists be verbalized to perform a tautology detection? Garbage collection algorithms to detect unimportant sentences in a paper draft?
VAs require multiple LLM calls by nature, which creates a performance bottleneck. An important avenue of future work includes leveraging efficient prefix caching, e.g., reusing the KV-cache of the prefix “Is X brighter than Y? X: $x_{1}$ Y:”, theoretically analyzing the “prefix-cache-friendliness” of various verbalized algorithms, and deriving the error bounds of VAs using past algorithmic literature on noisy oracles.
Appendix
Appendix A Sorting: Experimental Details
A.1 Bitonic Sorting Network
Bitonic sorting networks (Batcher’s Algorithm) (Batcher, 1968) allow us to optimize the sorting of a given list. The comparisons in a bitonic sorting network are not data-dependent, allowing for more feasible parallelization and speed proportional to depth $D$. Given an input list, a bitonic sorting network originally sorts both halves of the list, subsequently conducts an odd-even mergesort on both halves of the list and finally switches adjacent keys.
A sequence $a_{0},\ldots,a_{n-1}$ of $n$ elements is called bitonic when, for some $i$ (called an inflection point), $a_{0}\leq a_{1}...\leq a_{i}\geq a_{i+1}...\geq a_{n-1}$ (i.e., the first part up to $i$ is monotonically increasing and the second part is monotonically decreasing). Assume $n$ is a power of 2 and a sequence $a_{0},\ldots,a_{n-1}$ is bitonic with inflection point $i=\frac{n}{2}$. Then, we construct two sequences $s_{1}$ and $s_{2}$ with a following compare-and-exchange process:
$s_{1}$ $=\min(a_{0},a_{\frac{n}{2}}),\min(a_{1},a_{\frac{n}{2}+1}),...\min(a_{\frac{n}{2}-1},a_{n-1})$
$s_{2}$ $=\max(a_{0},a_{\frac{n}{2}}),\max(a_{1},a_{\frac{n}{2}+1}),...\max(a_{\frac{n}{2}-1},a_{n-1}).$
$s_{1}$ and $s_{2}$ are both bitonic sequences. Moreover, any element in $s_{2}$ are larger than any element in $s_{1}$, i.e., $\forall x\forall y$ $x\in s_{1}$, $x\in s_{2}$, $x<y$. The recursive compare-and-exchange process above is applied on $s_{1}$ and $s_{2}$ for $\log(n)$ times to sort both halves of the list.
For some input sizes, there are optimal sorting networks which require less comparisons or less parallel steps than Bitonic sorter does (Knuth, 1997). Bitonic sorting network can switch to such a network during the recursion, further increasing the performance.
A.2 Kendall-Tau Score
The Kendall Tau correlation coefficient Kendall (1945) is a non-parameteric statistic that measures how much order is preserved between two sets of rankings. Kendall-Tau score has several ways to compute ($\tau_{a},\tau_{b},\tau_{c}$) depending on the existence of ties or the nature of the observations. Here we describe $\tau_{b}$ we used in our paper.
Assume each observation consists of an input variable $X_{i}$ and an output variable $Y_{i}$. Consider two observations $(X_{i},Y_{i})$ and $(X_{j},Y_{j})$. These observations are called concordant if $X_{i}<X_{j}$ and $Y_{i}<Y_{j}$, or $X_{i}>X_{j}$ and $Y_{i}>Y_{j}$. Conversely, they are called discordant if $X_{i}<X_{j}$ and $Y_{i}>Y_{j}$, or $X_{i}>X_{j}$ and $Y_{i}<Y_{j}$. If $X_{i}=X_{j}$ and/or $Y_{i}=Y_{j}$, the observations are called tied.
Given $n$ observations $O$, enumerate all $n\choose 2$} pairs, which we denote as $N$. NN is then comprised of concordant pairs $P$, discordant pairs $Q$, pairs tied on the $X$ variable denoted as $T$, pairs tied on the $Y$ variable denoted as $U$, as well as pairs tied on both variables denoted as $B$. Kendall $\tau_{b}$ is then defined as:
$\tau_{b}=\frac{|P|-|Q|}{\sqrt{(|P|+|Q|+|T|)*(|P|+|Q|+|U|)}}\in[-1,1].$
A $\tau_{b}$ close to 1-1 indicates a negative correlation between two rankings and A $\tau_{b}$ close to $1$ indicates aligned rankings.
A.3 LLM Query Prompts
All methods uses a common shared criteria for comparison. In the Amazon dataset used in the study, the criteria prompt is:
criteria = ‘‘You will be given two product reviews, X and Y, written by various customers about the same product. Is X more favorable to the product than Y?’’
We used the following query prompts:
Constrained-decoding baseline:
f"Sort the following list in the increasing order of quality that is compared in the following comparison criteria: ’{criteria}’. That is, for a pair X and Y, if the answer to the criteria is ’yes’, X must appear after Y in the output list. "
I.i.d Scoring:
[…]
Autoregressive Scoring:
[…]
Verbalized Sorting:
f"{criteria} Answer with a single word, yes or no. X: ’{self.string}’ Y: ’{other.string}’"
Appendix B Clustering: Experimental Details
B.1 Jaccard distance
This implements a baseline clustering method using lexical similarity. Let $T_{i}$ represent the set of tokens from a string $s_{i}\in S$. For any two strings $s_{i}$ and $s_{j}$ with corresponding token sets $T_{i}$ and $T_{j}$, the Jaccard distance is computed as:
$J(s_{i},s_{j})=1-\frac{|T_{i}\cap T_{j}|}{|T_{i}\cup T_{j}|}$
where $|T_{i}\cap T_{j}|$ represents the cardinality of the intersection and $|T_{i}\cup T_{j}|$ represents the cardinality of the union.
B.2 t-STE
Van Der Maaten and Weinberger (2012) proposed a heavy-tailed kernel to measure local similarities between data points. In particular, we opt to use a Student-$t$ kernel with $\alpha$ degrees of freedom by defining:
$p_{ij\ell}=\frac{\left(1+\frac{\lVert x_{i}-x_{j}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}}{\left(1+\frac{\lVert x_{i}-x_{j}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}+\left(1+\frac{\lVert x_{i}-x_{\ell}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}}$
We refer to the resulting technique as t-Distributed STE (t-STE). Our implementation sets $\alpha=D-1$, where $D$ is the number of dimensions in the resulting embedding.
The formulation of the triplet probabilities is motivated by the success of unsupervised dimensionality reduction technique that also employ heavy-tailed similarity kernels (Maaten and Hinton, 2008).
B.3 LLM Query Prompts
Triplet comparison query uses the following shared prompt, where criteria is a dataset-specific prompt.
[…]
The dataset-specific prompt is as follows:
Medarxiv dataset:
criteria = f"You will be given three paper titles X, Y, and Z. Does the paper Z discuss a topic that is closer to the topic in X than to the topic in Y?"
IT Defect dataset:
criteria = f"You will be given three defect reports X, Y, and Z, sent by the customers of our GPFS filesystem storage system. They may contain stack traces of various programs (python, kernel drivers, ...). Similar repetitive lines are omitted as [...]. Considering the cause of the defects (i.e. ignoring timestamp/IP/MAC addresses), are X and Z more similar than Y and Z are? "
xkcdcolors-h dataset:
criteria = f"Which of the following two colors, X and Y, is more similar to the color Z? Focus on the hue in the HSL model first, then other aspects such as lightness and saturation. For example, Red and Pink are more similar than Red and Blue are. If you are not familiar with the given color name, e.g., a made-up color name like ’puke brown’, try to imagine what the color would look like, e.g., puke would generally have a lighter shade. "
xkcdcolors-l dataset:
criteria = f"Which of the following two colors, X and Y, is more similar to the color Z? Focus on the lightness in the HSL model first, then other aspects such as hue and saturation. For example, Red and Blue (both lightness = 50 If you are not familiar with the given color name, e.g., a made-up color name like ’puke brown’, try to imagine what the color would look like, e.g., puke would generally have a lighter shade. "
B.4 IT Defect Reports Dataset
Let $D$ denote the set of all defects. A subset $D^{\prime}\subseteq D$ was manually labeled as duplicates of several other defects in $D$. We selected 78 data points so that the final dataset contains 25 such clusters, each of size at least 2.
Since these labels are directional and sparse, we define the cluster by making the link undirected and computing the connected components, then treating a set of elements in each connected component as a cluster.
The following is a redacted/simplified defect records that appear in the dataset.
Defect 1
  "id": 347852,
  "description":
  "System login:
    EMS - X.XX.XXX.XXX root/cluster
    S6K w/ Falcon attached - lothal-qa6-[1 && 2] Code levels:
    S6k - 6230_B19
    FCM4 - 4_3_3
    Falcon - 5342 --> 5343
   Snap location:
    glogin:/u/DUMPS/R.347852/
   What I was doing:
    I was upgrading Falcon enclosure from 5342 to 5343.
   What I expected:
    ’XXXXfirmware’ will start and the Falcon will go through the upgrade process.
   What happened:
    ’mmchrimware’ completed but there are errors printed.
    The Falcon does appear to have the updated firmware.
    However, there are issues when trying to get a print
    from the Left module when running ’essenclosures’.
   Analysis:
    sg91:13:/dev/sg91::naa.500093d006585000/10910::[1.0.91.0]:/dev/sg91:XXX:5149-091:5342
    [10:48:31.258608008] XXXXXXFirmwareVer2 checking if enclosure needs to be updated
    [10:48:31.260567945] XXXXXXFirmwareVer2 firmware file found
    [10:48:31.262950549] XXXXXXFirmwareVer2 -> ...
    [10:48:31.350737778] XXXXXXFirmwareVer2 ...
    [10:48:31.362309567] XXXXXXFirmwareVer2 lothal-qa6-1-hs: [E] Error 1 updating firmware ...
    sg_ses failed: Transport error
  "duplicate_of": [],
Defect 2
This is marked as a duplicate of the previous defect.
  "id": 350018,
  "description": "Starting point:
    [root@s6krw3a ~]# XXXXfirmware --type storage-enclosure enclosure firmware ...
    enclosure 5149-091 78C100G 5342,5342 *5343 s6krw3a-hs enclosure 5149-F48 ...
    [root@s6krw3a ~]#
    Scenario:
    Storage enclosure fw update on 1 Falcon enclosure and 1 6k head unit. Build RC4
    [root@s6krw3a ~]# date; time XXXXfirmware --type storage-enclosure
    Thu May 29 05:26:19 PM EDT 2025
    Thu May 29 05:26:19 PM EDT 2025: XXXXfirmware: Processing node s6krw3a-hs....
    Thu May 29 05:26:23 PM EDT 2025: XXXXfirmware: Processing node s6krw3a-hs....
    s6krw3a-hs: update-directory .
    s6krw3a-hs: [I] Found storage-enclosure firmware update-id firmwareTable ...
    s6krw3a-hs: Updating enclosure firmware ESM_A.
    s6krw3a-hs: [E] Error 1 updating firmware for enclosure sg2.
    inquiry: transport: Host_status=0x0b [DID_SOFT_ERROR] /dev/sg2 doesn’t respond to a SCSI INQUIRY
    sg_ses failed: Transport error
    XXXXfirmware: Command failed. Examine previous error messages to determine cause.
    real 11m56.451s
    user 0m35.035s
    sys 2m7.818s
    [root@s6krw3a ~]#
    But apparently it still worked on canister A, but not on canister B
  "duplicate_of": [
    347852

### v3

@@
 Imagine receiving thousands of customer complaints and having to quickly sort them by how urgently they need support.
-Feeding such a large set of input strings blindly into an LLM is ineffective and impractical, both due to its limited context length and its limited ability to perform sorting tasks.
+Feeding such a large set of input strings blindly into a large language model (LLM) is ineffective and impractical, both due to its limited context length and its limited ability to perform sorting tasks.
 LLMs provide no guarantees on the correctness of their outputs for computational reasoning tasks, such as sorting, beyond anecdotal, empirical observations that they sometimes do.
+Despite the recent improvements, their behavior is fundamentally an approximation of formal reasoning, and thus contains hallucinations.
 In contrast, traditional disciplines of computer science often approach such a task using task decomposition.
@@
 These traditional algorithms, given an accurate oracle, guarantee the correctness of their outputs and the average/worst-case runtime.
+Such an algorithmic and analytical approach is often missing in LLM-based reasoning literature due to opaqueness of the reasoning mechanism in LLMs.
+More precisely, it often fails to distinguish the various separate challenges of reasoning tasks, i.e., (1) the ability to accurately perform atomic operations on ambiguous representations, (2) the ability to accurately execute a (not necessarily efficient) algorithm using those operations, (3) the choice of the algorithm to run and its theoretical properties (runtime/memory complexity, parallelism, robustness, soundness, completeness, etc.), and (4) the heuristic policy/hints for a branching algorithm.
+We argue that most reasoning tasks are informal representations of formal tasks for which an ideal classical algorithm with known complexity guarantees already exists in the literature, thus the effort to further improve (2,3) via additional training is wasteful and unproductive.
+For example, for the comparison sort, $O(\log n)$ algorithms are asymptotically optimal.
 In this work, we propose an abstract framework called “verbalized algorithms”.
-Given an informal representation of a formal task, a verbalized algorithm operates by replacing some basic atomic operations in a classical algorithm with LLM-driven text-based ones while retaining the high-level control flow of the initial algorithm, thereby largely retaining the theoretical guarantees of the original algorithm.
+Given an informal representation of a formal task, a verbalized algorithm operates by replacing some basic atomic operations in a classical algorithm with LLM-driven text-based ones while retaining the high-level control flow of the classical algorithm, thereby largely retaining the original theoretical guarantees.
 This has advantages over another common approach to similar tasks that we call the “formalization-based approach”.
@@
 Existing work based on this approach includes LLM-based generation of PDDL (McDermott, 2000) for classical planning (Xie, 2020; Guan et al., 2023; Li et al., 2024; Fine-Morris et al., 2024; Oswald et al., 2024), Prover9 input language for theorem proving (Olausson et al., 2023), or SMT solver input for optimization (Hao et al., 2024).
-The issue with this approach is twofold: First, it requires grammar-following capability for a complex target formal language.
+The issue with this approach is threefold: First, it requires grammar-following capability for a complex target formal language.
 Second, while the downstream solver requires a semantically and logically accurate input, the vagueness and informality in the natural language input often cause LLMs to generate inaccurate and inconsistent translations.
+Third, the assumptions made in the target formal language limits the scope of the tasks that can be solved by this approach.
+For example, a LLM-based planning agent that converts the task into PDDL cannot handle an informal task that contains probabilistic, temporal, or partially-observable nature.
+Verbalized algorithms, on the other hand, can handle them robustly via the general and informal common sense reasoning ability for atomic operations.
 2 Verbalized Algorithms
+Auto-regressive generative model is a probability distribution $p({\mathbf{x}})=p({\textnormal{x}}_{1},\ldots,{\textnormal{x}}_{N})$ where the random variable $\mathbf{x}$} represents a vector of integers, typically associated with a word or a sub-word token in natural language.
+Denoting its subsequence from index $i$ to $j$ as ${\mathbf{x}}_{i:j}$, Transformer-based language models (LMs) Vaswani et al.
+(2017) decompose it into each conditional distribution modeled by a large set of parameters $\theta$: $p({\mathbf{x}})=p_{\theta}({\mathbf{x}}_{N}|{\mathbf{x}}_{1:N-1})p({\mathbf{x}}_{1:N-1})=\ldots=\prod_{i=j}^{N-1}p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})p({\mathbf{x}}_{1:j})$, where the prefix ${\mathbf{x}}_{1:j}$ is typically called a prompt and each conditional distribution $p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})$ performs a next-token prediction.
+To exhibit a trait typically useful for user interaction (e.g.
+instruction following), LLMs go through a series of parameter-efficient fine-tuning (PEFT) steps after pretraining on a large corpus.
+For example, in Direct Preference Optimization (Rafailov et al., 2023, DPO), the input is a pair of model outputs in which human users have indicated their preference over another and the model is fine-tuned to follow that preference.
+For a prompt $x$ and an answer pair $y_{1}$ and $y_{2}$, denote by $y_{1}\succ y_{2}|x$ that $y_{1}$ is preferred over $y_{2}$.
+DPO learns a Bradley-Terry reward model (Bradley and Terry, 1952, BT-model) for this preference, i.e., $p(y_{1}\succ y_{2}|x)$.
+The Plackett-Luce reward model (Plackett, 1975; Luce, 1959, PL-model) extends BT to a multinominal case that represents a ranking over choices.
 While whether LLMs can reason or not is up for both philosophical and scientific debates, it is more widely accepted that they can solve some much simpler tasks with high fidelity, such as text-based classification and question-answering.
-Assuming those capabilities, we propose verbalized algorithms (VAs), an LLM-based reasoning strategy that limits the scope of the LLM’s usage to simple tasks.
+This is analogous to say that general BT (and PL) reward models baked in the LLM are sufficient to perform such simple tasks.
+Assuming those capabilities and thus relying entirely on the accuracy of the reward models, we propose verbalized algorithms (VAs), an LLM-based reasoning strategy that limits the scope of the LLMs to simple tasks.
 Definition 1 (Verbalized Algorithms).
@@
 A Verbalized Algorithm instantiates $A[\tau]$ with $\tau=\texttt{str}$ by implementing $f[\texttt{str}]$ as a query to an LLM whose output is constrained to $\left\{\text{``yes''},\text{``no''}\right\}$} which maps to $\left\{\top,\bot\right\}$}.
-VAs have three categories: Naive VAs, which treat the LLM’s answer as ground truth, Robust VAs, which treat it as a noisy oracle with a certain error rate and quantify the rate of error in the final output, and Probabilistic VAs, which use the probability/logits of LLM outputs to explicitly quantify the uncertainty of the result and return a distribution of outputs.
-In this workshop paper, we focus on the naive and the robust variants due to space, leaving the probabilistic variants for future work.
-2.1 Verbalized Sorting
+VAs keep the high-level control flow of $A[\tau]$ intact, while only replacing the oracle $f[\tau]$ with an LLM-based $f[\texttt{str}]$.
+This preserves the theoretical properties of the classical algorithm $A[\tau]$ if the LLM learned a perfect reward model that is equivalent to the ground truth.
+If the reward model is inaccurate, a non-classical robust algorithm that assumes noisy oracles can bound the error rate of the output using known theoretical results.
+VAs can also be combined with other specialized algorithms, such as concurrent algorithms.
+While our definition is focused on the BT reward model, it naturally extends to PL rewards / categorical oracles.
+VAs have three categories: Naive VAs, which treat the LLM’s answer as ground truth, Robust VAs, which treat it as a noisy oracle with a certain error rate and quantify the rate of error in the final output, and Probabilistic VAs, which use the probability/logits of LLM outputs to improve the accuracy, to quantify the uncertainty of the result, or to return a distribution of outputs.
+3 Verbalized Sorting
+Table 1: Average number of sorting invariance violations by the constraint decoding baseline.
+“Duplicates” indicates the number of duplicates found in the output, and “missing” indicates the number of input elements that are missing in the output, for 25 input strings.
+Models (Qwen3) duplicates missing 1.7B 1.39 9.63 32B-AWQ 0.49 2.75
+Table 2: Average Kendall-Tau scores over 5 seeds. Best scores in bold. The score of 1, 0, and -1 indicates perfect rank correlations, no correlations, and perfect anti-correlations, respectively.
+Models Scoring Naive VA Robust ($K=3$) Symmetrized (Qwen3) Baseline i.i.d.
+AR Bitonic Powersort Bitonic Powersort Bitonic Powersort 1.7B 0.00 0.18 0.05 0.30 0.33 0.33 0.44 0.33 0.48 32B-AWQ 0.30 0.57 0.39 0.39 0.56 0.39 0.58 0.41 0.60
+Table 3: Average total elapsed time in seconds for processing 100 sorting tasks (each with 25 strings). Bitonic sorting network also shows the relative speedup over the sequential Powersort.
+Models Scoring Naive VA Robust ($K=3$) (Qwen3) Baseline i.i.d.
+AR Power.
+Bitonic Power Bitonic Bitonic+16 1.7B 9279 636 194 5196 2476 (x2.1) 7279 5304 (x1.4) 4728 (x1.5) 32B-AWQ 6798 1406 661 18820 8380 (x2.3) 25302 20461 (x1.2) 17169 (x1.5)
+Table 4: Average total elapsed time of naive Powersort in seconds, comparing VLLM without kv-cache sharing, with kv-cache sharing via PagedAttention, and SGLang with RadixAttention.
+Models VLLM (PagedAttention) SGLang (Qwen3) disabled enabled (RadixAttention) 1.7B 5196 5403 (x.96) 3812 (x1.4) 32B-AWQ 18820 17421 (x1.1) 14424 (x1.3)
+Table 5: Clustering scores across methods and benchmarks. The top 2 results are highlighted in bold.
+method model $D=$ $k=$ Medarxiv IT Defects xkcdcolors-h xkcdcolors-l Jaccard 8 0.556 0.767 0.601 0.472 ST Qwen3-Emb…d-8B 4096 0.827 0.936 0.718 0.531 triplet Qwen3-4B 8 100 0.831 0.831 0.823 0.704 triplet Qwen3-4B 8 200 0.858 0.895 0.823 0.691 triplet Qwen3-32B-AWQ 8 100 0.842 0.908 0.851 0.748 triplet Qwen3-32B-AWQ 8 200 0.858 0.939 0.850 0.742
 Verbalized Sorting sorts a list of natural language strings $X$ using LLM’s question answering capability as an atomic comparison oracle.
@@
 The model output is mapped to $\top/\bot$ if the sampled output is “yes” / “no”, respectively.
+This task is important in the RAG applications and also called the LLM-based ranking approach.
 We evaluated the accuracy of several non-VA and VA approaches.
@@
 Due to the lack of a computationally efficient way to constrain the list to (1) be free of duplicates and (2) represent the same set of strings as the input, we (1) remove the duplicates and (2) append missing elements to the end of the list.
+The approach corresponds to listwise ranking approach Ma et al.
+(2023) in the RAG literature.
 Scoring-based approaches ask the model to assign a score between 1.0 and 5.0 (down to one decimal) to each sentence using regex-constrained decoding, then sort the sentences using the scores.
@@
 We tested two variants: I.i.d Scoring scores $x_{i}\in X$ independently, and Autoregressive Scoring scores all strings in a single query, allowing the model to see the scores of other elements.
+The approach corresponds to pointwise ranking approach Sun et al.
+(2023); Sachan et al.
+(2022) in the RAG literature.
+VA approaches correspond to pairwise ranking approaches Qin et al.
+(2024) in the RAG literature, but our contribution focuses on leveraging the theoretical properties of classical algorithms, which has been undermined in the literature.
 VA Powersort overrides Python’s __gt__ operator with an LLM call and applies the built-in sorted function, which uses Powersort (Munro and Wild, 2018) that runs $O(n\log n)$ comparisons worst-case, the best runtime complexity achievable by comparison sort.
@@
 It performs $O(n(\log n)^{2})$ comparisons, but runs in $O((\log n)^{2})$ time due to parallelism.
-VA Robust Powersort/Bitonic Sorting Network makes individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting.
+As an additional optimization, we use the optimal sorting network (Knuth, 1997; Bundala and Závodnỳ, 2014) when the input size is 16 during its recursion.
+Robust VAs (Powersort/Bitonic) make individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting.
 For an odd $K$, if the oracle is correct with probability $p$, the voting result is correct with at least probability $1-e^{-K(2p-1)^{2}/2}$ (Emamjomeh-Zadeh and Kempe, 2018) due to Hoeffding bound (Hoeffding, 1963).
+Finally, Probabilistic VAs (Powersort/Bitonic) instead rely on comparing the probabilities / logits over the yes/no answers to perform each comparison query.
+While the naive probabilistic VA is equivalent to the naive VA with greedy sampling, we additionally have Symmetrized VA to address two biases in LLM: Positional bias and psycophancy (the tendency to reply “yes”) .
+Let $p_{12}$ and $p_{21}$ be next-token probabilities for answering “yes” with the original prompt and a prompt whose $x_{1}$ and $x_{2}$ are swapped.
+Due to the bias, $p_{12}\not=1-p_{21}$.
+Symmetrized VA tests $\frac{p_{12}+(1-p_{21})}{2}>\frac{(1-p_{12})+p_{21}}{2}$ as a comparison oracle.
+Evaluation
 We evaluated the approaches above on Amazon reviews benchmark (Hou et al., 2024).
@@
 We used Qwen3-1.7B and Qwen3-32B-AWQ models (Team, 2025) with thinking enabled.
-First, Table 1 (1) shows how often the constraint-decoding baseline violates the sorting invariance, i.e., how many elements are duplicated or missing in the output on average.
+First, Table 1 shows how often the constraint-decoding baseline violates the sorting invariance, i.e., how many elements are duplicated or missing in the output on average.
 While these violations become rarer as the model gets larger, they do not disappear.
 This highlights the lack of theoretical guarantees in LLM-based reasoning, which does not occur in VAs and formalization-based approaches.
-Next, Table 1 (2) shows the average Kendall-Tau scores.
-Constraint-decoding baseline performed the worst overall, and while the scoring approaches can work with a large 32B model, their performance dwindles when the model is smaller.
+Next, Table 2 shows the average Kendall-Tau scores.
+Constraint-decoding baseline performed the worst overall, and while the scoring approaches can work with the large 32B model, their performance dwindles when the model is smaller.
 Robust VA with 32B achieved the best results.
-The strength of VA is most pronounced when using the small 1.7B model, where other approaches failed.
+The strength of VA is most pronounced with the small 1.7B model, where other approaches failed.
 Notably, VAs with a 1.7B model outperformed the baseline with a 32B model, clearly showing the advantage of using the sorting algorithms developed in the literature.
 Finally, Table 3 compares the runtime of each approach.
-VA Bitonic Sorting Network improves over the VA Powersort, and both VAs can be faster than the constrained-decoding baseline, even with robustification (x3 parallel queries).
-This shows that VAs can directly benefit from the theoretical results in the parallel sorting literature, in contrast to the recent LLM reasoning literature that tends to detach from existing discrete algorithm studies.
+VA Bitonic Sorting Network improves over the VA Powersort, and both VAs can be faster than the constrained-decoding baseline, even with robustification (x3 parallel queries) or symmetrization (x2 parallel queries).
+This shows that VAs can directly benefit from the theoretical results in the parallel sorting literature.
 Note that the parallel speedup in bitonic sort does not reach the theoretical limit (x11.4; 171 comparators and 15 steps for 25 inputs) because the maximum batch size is limited by the GPU; multi-GPU experiments with data-parallelism would improve this.
-Table 1: (1) Average number of sorting invariance violations by the constraint decoding baseline.
-“Duplicates” indicates the number of duplicates found in the output, and “missing” indicates the number of input elements that are missing in the output, for 25 input strings.
-(2) Average Kendall-Tau scores.
-Top-3 are highlighted in bold.
-The score of 1, 0, and -1 indicates perfect rank correlations, no correlations, and perfect anti-correlations, respectively.
-(1) Invariance (2) Accuracy (Kendall-Tau scores) Models Violations Constraint Scoring Naive VA Robust VA ($K=3$) (Qwen3) duplicates missing Decoding i.i.d.
-AR Bitonic Powersort Bitonic Powersort 1.7B 1.39 9.63 0.00 0.18 0.05 0.30 0.33 0.33 0.44 32B-AWQ 0.49 2.75 0.30 0.57 0.39 0.39 0.56 0.39 0.58
-Table 2: Average total elapsed time in seconds for processing all sorting queries. Bitonic sorting network also shows the relative speedup over the sequential sort (Powersort).
-Models Scoring Naive VA Robust VA ($K=3$) (Qwen3) Baseline i.i.d.
-AR Bitonic Powersort Bitonic Powersort 1.7B 7468 630 205 2338 (x2.12) 4975 5060 (x1.45) 7362 32B-AWQ 7280 1430 626 8556 (x2.18) 18679 20350 (x1.30) 26471
-2.2 Verbalized Clustering
+The Effect of Prefix Caching
+Each query in VA sort requires a single next-token prediction.
+In Transformer architecture, it requires $O(t^{2})$ computations for $t$ input tokens.
+However, State-of-the-Art inference engines such as vllm Kwon et al.
+(2023) and SGLang Zheng et al.
+(2024) cache the prefix kv-cache over multiple queries.
+This is particularly helpful in reducing the inference cost of VA-style queries majority of which share the prefix.
+A comparison query in VA sorting consists of a prompt, $x_{1}$, and $x_{2}$.
+Assume all strings being compared has a fixed length $L=|x_{1}|=|x_{2}|$, and that the length of the fixed prefix is $L_{p}$.
+In the exhaustive pairwise comparison among $n$ strings ($O(n^{2})$ comparisons, equivalent to insertion sort etc.), a naive inference engine requires $O((L_{p}+2L)^{2})$ runtime per each yes/no query, thus $O(n^{2}(L_{p}+2L)^{2})$ runtime in total, without parallelism.
+Assume that an input consists of $N+M$ tokens (prefix and suffix), and the kv-cache of $N$ prefix tokens is already available.
+Each $N+t$-th new input token attends to N+(t1)N+(t-1) existing tokens, which makes $N+M+1$-th token require $\sum_{t=1}^{M}(N+(t-1))=O(NM+M^{2})$ runtime.
+With this in mind, assuming that the fixed prompt already has a kv-cache, the runtime of each comparison query reduces to $O(2L_{p}L+4L^{2})$, totalling $O(L_{p}^{2}+n^{2}(2L_{p}L+4L^{2}))$ runtime.
+Moreover, since $x_{1}$ is repeated $n$ times over $n^{2}$ comparisons, it further reduces to $O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}((L_{p}+L)L+L^{2}))=O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}(L_{p}L+2L^{2}))$.
+This is significant because the prompt tends to be much longer than each string.
+Table 4 demonstrates the empirical speedup due to kv-cache in two State-of-the-Art inference engines. RadixAttention in SGLang is particularly helpful for VAs whose queries share large prefixes.
+4 Verbalized Clustering
 Verbalized Clustering attempts to cluster a set of strings $S$ into a family of subsets $\mathcal{S}$ based on their similarity.
@@
 The score is defined as:
-score $=\textstyle\quad\frac{1}{|S|}\sum_{s\in S}\frac{1}{|C(s)|-1}\sum_{s^{\prime}\in C(s)\setminus{\left\{s\right\}}}1-\frac{\textsc{rank}(s,s^{\prime})}{|S|}\in[0,1]$ (1)
+score $=\frac{1}{|S|}\sum_{s\in S}\frac{1}{|C(s)|-1}\sum_{s^{\prime}\in C(s)\setminus{\left\{s\right\}}}1-\frac{\textsc{rank}(s,s^{\prime})}{|S|}\in[0,1]$ (1)
 Table 5 shows the results.
@@
 We highlight its strong generalization in OOD tasks where other approaches failed, due to its strong algorithmic bias combined with flexible LLM queries.
-Table 3: Clustering scores across methods and benchmarks. The top 2 results are highlighted in bold.
-method model $D=$ $k=$ Medarxiv IT Defects xkcdcolors-h xkcdcolors-l Jaccard 8 0.556 0.767 0.601 0.472 ST Qwen3-Emb…d-8B 4096 0.827 0.936 0.718 0.531 triplet Qwen3-4B 8 100 0.831 0.831 0.823 0.704 triplet Qwen3-4B 8 200 0.858 0.895 0.823 0.691 triplet Qwen3-32B-AWQ 8 100 0.842 0.908 0.851 0.748 triplet Qwen3-32B-AWQ 8 200 0.858 0.939 0.850 0.742
-3 Conclusion, Related Work, Discussion and Future Work
+5 Related Work
+Despite significant recent findings in reasoning models, which combine RL-based self-training with “think tags” (Guo et al., 2025), they rarely explicitly perform systematic backtracking and search.
+Whether such systems can reason or not has been hotly debated in the context of LLM-based planning (Valmeekam et al., 2023).
+Previous work on solving planning tasks using LLM (Liu et al., 2023; Hirsch et al., 2024) demonstrates varying degrees of success and limitations (e.g., failure in mystery due to sensitivity to the vocabulary).
+Formalization approach to planning is also called a parse-and-solve model (Collins et al., 2022) as opposed to a LLM-as-Planner model that naively generates a plan with an LLM.
+Thought-of-Search (Katz et al., 2024) simulates a systematic search algorithm, which could be seen as an instance of VAs.
+To our knowledge, there is no work that simulates traditional graph search algorithms (e.g., A* (Hart et al., 1968), GBFS (Bonet and Geffner, 2001)) in a VA manner, i.e., by replacing all of its atomic operations (e.g., successor generation, heap-based priority queues, hash computation for duplicate detection) with LLM queries, although some work replace some atomic operations, especially successor generations (Katz et al., 2024).
+A number of recent papers in the Algorithm with Predictions (Mitzenmacher and Vassilvitskii, 2022, ALPS) community achieved a significant success in augmenting algorithms by data-driven oracles.
+The particular strength of this line of work is the rigorous theoretical analysis of the error rate.
+VAs is a form of ALPS specialized in LLM-based oracles, thus combining ideas from ALPS community with VAs is a promising direction for future work.
+Combining data-driven heuristics with existing algorithms in the classical AI tasks is also a common practice in the recent literature (e.g., classical planning Núñez-Molina et al.
+(2024), boolean SAT Duan et al.
+(2020), logistics tasks Xin et al.
+(2021)).
+6 Conclusion, Discussion and Future Work
 We presented Verbalized Algorithms (VAs), a disciplined paradigm for performing algorithmic reasoning with LLMs.
@@
 It also leverages their established theoretical results to improve the runtime of such reasoning.
-Despite significant recent findings in reasoning models, which combine RL-based self-training with “think tags” (Guo et al., 2025), they rarely explicitly perform systematic backtracking and search.
-Whether such systems can reason or not has been hotly debated in the context of LLM-based planning (Valmeekam et al., 2023).
-Previous work on solving planning tasks using LLM (Liu et al., 2023; Hirsch et al., 2024) demonstrates varying degrees of success and limitations (e.g., failure in mystery due to sensitivity to the vocabulary).
-Thought-of-Search (Katz et al., 2024) simulates a systematic search algorithm, which could be seen as an instance of VAs.
 While this paper explored only a limited number of applications due to space, we are already pursuing several exciting verbalizations, including submodular maximization and pareto-front computation for diverse / non-dominated document retrieval, DPLL for combinatorial reasoning, and median/top-$k$ selection, etc.
@@
 Garbage collection algorithms to detect unimportant sentences in a paper draft?
-VAs require multiple LLM calls by nature, which creates a performance bottleneck.
-An important avenue of future work includes leveraging efficient prefix caching, e.g., reusing the KV-cache of the prefix “Is X brighter than Y?
-X: $x_{1}$ Y:”, theoretically analyzing the “prefix-cache-friendliness” of various verbalized algorithms, and deriving the error bounds of VAs using past algorithmic literature on noisy oracles.
 Appendix
@@
 We selected 78 data points so that the final dataset contains 25 such clusters, each of size at least 2.
-Since these labels are directional and sparse, we define the cluster by making the link undirected and computing the connected components, then treating a set of elements in each connected component as a cluster.
-The following is a redacted/simplified defect records that appear in the dataset.
-Defect 1
+Example Datapoint (redacted/simplified)
+The following is an example of a redacted/simplified defect record that appear in the dataset.
   "id": 347852,
@@
   "duplicate_of": [],
-Defect 2
-This is marked as a duplicate of the previous defect.
-  "id": 350018,
-  "description": "Starting point:
-    [root@s6krw3a ~]# XXXXfirmware --type storage-enclosure enclosure firmware ...
-    enclosure 5149-091 78C100G 5342,5342 *5343 s6krw3a-hs enclosure 5149-F48 ...
-    [root@s6krw3a ~]#
-    Scenario:
-    Storage enclosure fw update on 1 Falcon enclosure and 1 6k head unit. Build RC4
-    [root@s6krw3a ~]# date; time XXXXfirmware --type storage-enclosure
-    Thu May 29 05:26:19 PM EDT 2025
-    Thu May 29 05:26:19 PM EDT 2025: XXXXfirmware: Processing node s6krw3a-hs....
-    Thu May 29 05:26:23 PM EDT 2025: XXXXfirmware: Processing node s6krw3a-hs....
-    s6krw3a-hs: update-directory .
-    s6krw3a-hs: [I] Found storage-enclosure firmware update-id firmwareTable ...
-    s6krw3a-hs: Updating enclosure firmware ESM_A.
-    s6krw3a-hs: [E] Error 1 updating firmware for enclosure sg2.
-    inquiry: transport: Host_status=0x0b [DID_SOFT_ERROR] /dev/sg2 doesn’t respond to a SCSI INQUIRY
-    sg_ses failed: Transport error
-    XXXXfirmware: Command failed. Examine previous error messages to determine cause.
-    real 11m56.451s
-    user 0m35.035s
-    sys 2m7.818s
-    [root@s6krw3a ~]#
-    But apparently it still worked on canister A, but not on canister B
-  "duplicate_of": [
-    347852

### v4

@@
-Verbalized Algorithms
+Verbalized Algorithms: Zero-shot Classical Algorithmic Reasoning for Correctness and Runtime Guarantees
 Abstract
-Instead of querying LLMs in a one-shot manner and hoping to get the right answer for a reasoning task, we propose a paradigm we call verbalized algorithms (VAs), which leverage classical algorithms with established theoretical understanding.
-VAs decompose a task into simple elementary operations on natural language strings that they should be able to answer reliably, and limit the scope of LLMs to only those simple tasks.
+Instead of querying LLMs in a one-shot manner and hoping to get the right answer for a reasoning task, we propose a paradigm we call verbalized algorithms (VAs), which combines LLMs with classical algorithms with established theoretical guarantees.
+VAs decompose a task into simple elementary operations on natural language strings that LLMs are able to answer reliably, and limit the scope of LLMs to those simple tasks.
 For example, for sorting a series of natural language strings, verbalized sorting uses an LLM as a binary comparison oracle in a known and well-analyzed sorting algorithm (e.g., bitonic sorting network).
-We demonstrate the effectiveness of this approach on sorting and clustering tasks.
+Although this is already known as pairwise ranking in the literature, we additionally demonstrate the effectiveness of verbalized maximum, verbalized clustering, and verbalized submodular maximization for numerical reasoning, topic clustering and multi-hop Q&A RAG task, which guarantees $O(n)$ runtime, $O(n\log n)$ runtime, and 1/(1e)1/(1-e) optimality, respectively.
+Clustering and submodular maximization outperformed or improved the nearest neighbor search using state-of-the-art embedding models.
+Large Language Model, Reasoning
+[…]
 1 Introduction
@@
 In contrast, traditional disciplines of computer science often approach such a task using task decomposition.
-For example, a merge sort decomposes the entire sorting task into several sorting tasks over individual sublists and combines the results, where, at the limit, there is an elementary binary comparison operator that compares individual pairs of objects in the list.
-These traditional algorithms, given an accurate oracle, guarantee the correctness of their outputs and the average/worst-case runtime.
+For example, a merge sort decomposes the entire sorting task into smaller sorting tasks and combines the results, where, at the limit, there is an elementary binary comparison that compares an individual pair.
+These traditional algorithms, given an accurate oracle, guarantee the correctness of their outputs and the average/worst-case runtime/memory.
 Such an algorithmic and analytical approach is often missing in LLM-based reasoning literature due to opaqueness of the reasoning mechanism in LLMs.
-More precisely, it often fails to distinguish the various separate challenges of reasoning tasks, i.e., (1) the ability to accurately perform atomic operations on ambiguous representations, (2) the ability to accurately execute a (not necessarily efficient) algorithm using those operations, (3) the choice of the algorithm to run and its theoretical properties (runtime/memory complexity, parallelism, robustness, soundness, completeness, etc.), and (4) the heuristic policy/hints for a branching algorithm.
-We argue that most reasoning tasks are informal representations of formal tasks for which an ideal classical algorithm with known complexity guarantees already exists in the literature, thus the effort to further improve (2,3) via additional training is wasteful and unproductive.
-For example, for the comparison sort, $O(\log n)$ algorithms are asymptotically optimal.
+More precisely, it often fails to distinguish various separate challenges associated with reasoning, such as (1) the ability to accurately perform atomic operations on ambiguous representations, (2) the ability to accurately execute a sound but not necessarily efficient algorithm using those operations, (3) the theoretical efficiency of the algorithm (runtime/memory complexity, parallelism, robustness, soundness, completeness, etc.), and (4) the heuristic policy/hints for a branching algorithm.
+We argue that most reasoning tasks are informal representations of formal tasks for which an asymptotically optimal classical algorithm already exists in the literature, thus the effort to further improve (2,3) via additional training is wasteful and unproductive.
+For example, for comparison sort, $O(n\log n)$ algorithms are asymptotically optimal.
 In this work, we propose an abstract framework called “verbalized algorithms”.
-Given an informal representation of a formal task, a verbalized algorithm operates by replacing some basic atomic operations in a classical algorithm with LLM-driven text-based ones while retaining the high-level control flow of the classical algorithm, thereby largely retaining the original theoretical guarantees.
+Given an informal representation of a formal task, a verbalized algorithm replaces some atomic operations in a classical algorithm with LLM-driven text-based oracles while retaining the high-level control flow, thereby largely retaining the theoretical guarantees.
 This has advantages over another common approach to similar tasks that we call the “formalization-based approach”.
@@
 For example, a LLM-based planning agent that converts the task into PDDL cannot handle an informal task that contains probabilistic, temporal, or partially-observable nature.
-Verbalized algorithms, on the other hand, can handle them robustly via the general and informal common sense reasoning ability for atomic operations.
+Verbalized algorithms, on the other hand, can handle them robustly by applying a general and informal common sense reasoning to atomic operations.
+This paper is organized as follows.
+We first formalize verbalize algorithms.
+We then perform four case studies with increasing difficulty and non-triviality: Maximum, sorting, clustering, and submodular maximization.
+Each section contains a task description, alternative algorithms considered, and an empirical evaluation.
+All experimental results are averages of 5 random seeds.
 2 Verbalized Algorithms
 Auto-regressive generative model is a probability distribution $p({\mathbf{x}})=p({\textnormal{x}}_{1},\ldots,{\textnormal{x}}_{N})$ where the random variable $\mathbf{x}$} represents a vector of integers, typically associated with a word or a sub-word token in natural language.
-Denoting its subsequence from index $i$ to $j$ as ${\mathbf{x}}_{i:j}$, Transformer-based language models (LMs) Vaswani et al.
-(2017) decompose it into each conditional distribution modeled by a large set of parameters $\theta$: $p({\mathbf{x}})=p_{\theta}({\mathbf{x}}_{N}|{\mathbf{x}}_{1:N-1})p({\mathbf{x}}_{1:N-1})=\ldots=\prod_{i=j}^{N-1}p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})p({\mathbf{x}}_{1:j})$, where the prefix ${\mathbf{x}}_{1:j}$ is typically called a prompt and each conditional distribution $p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})$ performs a next-token prediction.
+Denoting its subsequence from index $i$ to $j$ as ${\mathbf{x}}_{i:j}$, Transformer-based language models (LMs) (Vaswani et al., 2017) decompose it into conditional distributions modeled by a large set of parameters $\theta$: $p({\mathbf{x}})=p_{\theta}({\mathbf{x}}_{N}|{\mathbf{x}}_{1:N-1})p({\mathbf{x}}_{1:N-1})=\ldots=\prod_{i=j}^{N-1}p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})p({\mathbf{x}}_{1:j})$, where the prefix ${\mathbf{x}}_{1:j}$ is typically called a prompt and each conditional distribution $p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})$ performs a next-token prediction.
 To exhibit a trait typically useful for user interaction (e.g.
@@
 For a prompt $x$ and an answer pair $y_{1}$ and $y_{2}$, denote by $y_{1}\succ y_{2}|x$ that $y_{1}$ is preferred over $y_{2}$.
-DPO learns a Bradley-Terry reward model (Bradley and Terry, 1952, BT-model) for this preference, i.e., $p(y_{1}\succ y_{2}|x)$.
+DPO learns a Bradley-Terry reward model (Bradley & Terry, 1952, BT-model) for this preference, i.e., $p(y_{1}\succ y_{2}|x)$.
 The Plackett-Luce reward model (Plackett, 1975; Luce, 1959, PL-model) extends BT to a multinominal case that represents a ranking over choices.
@@
 VAs have three categories: Naive VAs, which treat the LLM’s answer as ground truth, Robust VAs, which treat it as a noisy oracle with a certain error rate and quantify the rate of error in the final output, and Probabilistic VAs, which use the probability/logits of LLM outputs to improve the accuracy, to quantify the uncertainty of the result, or to return a distribution of outputs.
+3 Verbalized Maximum
+As a simplest algorithmic task to evaluate the reasoning accuracy of LLMs, we first investigate its ability to find the maximum number in a list $X$ of $n$ randomly selected integers between 1 and 10000.
+Due to the lack of principled algorithm for the LLM to follow, a naive LLM often fails even at this simple task despite with the help of constrained decoding, especially when the list gets longer.
+We compare two approaches: Constraint Decoding Baseline queries LLMs with a string “What is the largest number in [comma-separated numbers in XX]?”.
+We use xgrammar constraint decoding engine to restrict the space of possible outputs to one of the elements in $X$.
+Verbalized maximum overrides Python’s __gt__ operator with an LLM call and applies the built-in max function, which sequentially reduces the list with $O(n)$ comparisons.
+Each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query “Is X larger than Y?
+X: $x_{1}$ Y: $x_{2}$”, where the next output token is restricted to “yes” and “no” with constrained decoding.
+The model output is mapped to $\top/\bot$ if the sampled output is “yes” / “no”, respectively.
+We evaluated the algorithms above across 20 randomly sampled lists for each of $n\in{\left\{10,100\right\}}$ with various Qwen3 models (1.7B, 4B, 32B) with and without thinking enabled.
+We measured the mean absolute error (MAE) between the ground-truth maximum and the returned maximum.
+Note that the returned answer is guaranteed to be an element of $X$ in both algorithms.
+The result is shown in Table 1.
+MAE column suggests that the baseline with a small 1.7B model struggles even on a small input $n=10$ and with the help of thinking.
+On larger inputs $(n=100)$, the accuracy degrades further on various model sizes, while verbalized maximum maintains a small, often zero error.
+Table 1: Mean absolute error between the ground truth maximum and the LLM’s answer, and the runtime.
+Models MAE($n=10$) MAE($n=100$) (Qwen3) baseline verbalized baseline verbalized 1.7B 500.14 0 1418.9 9.45 1.7B+think 14.18 0 1071.6 38.95 4B 0 0 34.25 0 4B+think 0 0 49.85 0 32B 0 0 48.6 0 32B+think 0 0 11.45 0
 4 Verbalized Sorting
+Verbalized Sorting sorts a list of natural language strings $X$ using LLM’s question answering capability as an atomic comparison oracle.
+Similar to the verbalized maximum, each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query, although the exact prompt varies depending on the target application (see appendix Sec.
+A for details).
+This task is important in the RAG applications and also called the LLM-based ranking approach.
+We evaluated the accuracy of several non-VA and VA approaches.
+Constraint Decoding Baseline encodes the list $X$ in a markdown format and asks the model to return a sorted list using structured-decoding (Willard & Louf, 2023) to constrain the output to be a list whose elements are guaranteed to be one of the input strings.
+Due to the lack of a computationally efficient way to constrain the list to (1) be free of duplicates and (2) represent the same set of strings as the input, we (1) remove the duplicates and (2) append missing elements to the end of the list.
+The approach corresponds to listwise ranking approach (Ma et al., 2023) in the RAG literature.
+Scoring-based approaches ask the model to assign a score between 1.0 and 5.0 (down to one decimal) to each sentence using regex-constrained decoding, then sort the sentences using the scores.
+It is a form of formalization-based approach (the list of scores represents the formal problem).
+We tested two variants: I.i.d Scoring scores $x_{i}\in X$ independently, and Autoregressive Scoring scores all strings in a single query, allowing the model to see the scores of other elements.
+The approach corresponds to pointwise ranking approach (Sun et al., 2023; Sachan et al., 2022) in the RAG literature.
+VA approaches correspond to pairwise ranking approaches (Qin et al., 2024) in the RAG literature, but our contribution focuses on leveraging the theoretical properties of classical algorithms, which has been undermined in the literature.
+VA Powersort overrides Python’s __gt__ operator with an LLM call and applies the built-in sorted function, which uses Powersort (Munro & Wild, 2018) that runs $O(n\log n)$ comparisons worst-case, the best runtime complexity achievable by comparison sort.
+VA Bitonic sorting network uses the Bitonic sorting network (Batcher, 1968) from parallel sorting literature to improve the throughput.
+It performs $O(n(\log n)^{2})$ comparisons, but runs in $O((\log n)^{2})$ time due to parallelism.
+As an additional optimization, we use the optimal sorting network (Knuth, 1997; Bundala & Závodnỳ, 2014) when the input size is 16 during its recursion.
+Robust VAs (Powersort/Bitonic) make individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting.
+For an odd $K$, if the oracle is correct with probability $p$, the voting result is correct with at least probability $1-e^{-K(2p-1)^{2}/2}$ (Emamjomeh-Zadeh & Kempe, 2018) due to Hoeffding bound (Hoeffding, 1963).
+Finally, Probabilistic VAs (Powersort/Bitonic) instead rely on comparing the probabilities / logits over the yes/no answers to perform each comparison query.
+While the naive probabilistic VA is equivalent to the naive VA with greedy sampling, we additionally have Symmetrized VA to address two biases in LLM: Positional bias (Liu et al., 2024) and psycophancy (the tendency to reply “yes”) (Casper et al., ).
+Let $p_{12}$ and $p_{21}$ be next-token probabilities for answering “yes” with the original prompt and a prompt whose $x_{1}$ and $x_{2}$ are swapped.
+Due to the bias, $p_{12}\not=1-p_{21}$.
+Symmetrized VA tests $\frac{p_{12}+(1-p_{21})}{2}>\frac{(1-p_{12})+p_{21}}{2}$ as a comparison oracle.
 Table 2: Average number of sorting invariance violations by the constraint decoding baseline.
 “Duplicates” indicates the number of duplicates found in the output, and “missing” indicates the number of input elements that are missing in the output, for 25 input strings.
-Models (Qwen3) duplicates missing 1.7B 1.39 9.63 32B-AWQ 0.49 2.75
-Table 2: Average Kendall-Tau scores over 5 seeds. Best scores in bold. The score of 1, 0, and -1 indicates perfect rank correlations, no correlations, and perfect anti-correlations, respectively.
+Models duplicates missing Qwen3-1.7B 1.39 9.63 Qwen3-32B-AWQ 0.49 2.75
+Table 3: Average Kendall-Tau scores over 5 seeds. Larger the better, best scores in bold.
 Models Scoring Naive VA Robust ($K=3$) Symmetrized (Qwen3) Baseline i.i.d.
@@
 Models VLLM (PagedAttention) SGLang (Qwen3) disabled enabled (RadixAttention) 1.7B 5196 5403 (x.96) 3812 (x1.4) 32B-AWQ 18820 17421 (x1.1) 14424 (x1.3)
-Table 5: Clustering scores across methods and benchmarks. The top 2 results are highlighted in bold.
-method model $D=$ $k=$ Medarxiv IT Defects xkcdcolors-h xkcdcolors-l Jaccard 8 0.556 0.767 0.601 0.472 ST Qwen3-Emb…d-8B 4096 0.827 0.936 0.718 0.531 triplet Qwen3-4B 8 100 0.831 0.831 0.823 0.704 triplet Qwen3-4B 8 200 0.858 0.895 0.823 0.691 triplet Qwen3-32B-AWQ 8 100 0.842 0.908 0.851 0.748 triplet Qwen3-32B-AWQ 8 200 0.858 0.939 0.850 0.742
-Verbalized Sorting sorts a list of natural language strings $X$ using LLM’s question answering capability as an atomic comparison oracle.
-Each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query such as “Is X better than Y?
-X: $x_{1}$ Y: $x_{2}$”, where the next output token is restricted to “yes” and “no”.
-The model output is mapped to $\top/\bot$ if the sampled output is “yes” / “no”, respectively.
-This task is important in the RAG applications and also called the LLM-based ranking approach.
-We evaluated the accuracy of several non-VA and VA approaches.
-Method, dataset, and prompt details are available in the appendix Sec.
-A.
-All results are averages of 5 random seeds.
-Constraint Decoding Baseline encodes the list $X$ in a markdown format and asks the model to return a sorted list using structured-decoding (Willard and Louf, 2023) to constrain the output to be a list whose elements are guaranteed to be one of the input strings.
-Due to the lack of a computationally efficient way to constrain the list to (1) be free of duplicates and (2) represent the same set of strings as the input, we (1) remove the duplicates and (2) append missing elements to the end of the list.
-The approach corresponds to listwise ranking approach Ma et al.
-(2023) in the RAG literature.
-Scoring-based approaches ask the model to assign a score between 1.0 and 5.0 (down to one decimal) to each sentence using regex-constrained decoding, then sort the sentences using the scores.
-It is a form of formalization-based approach (the list of scores represents the formal problem).
-We tested two variants: I.i.d Scoring scores $x_{i}\in X$ independently, and Autoregressive Scoring scores all strings in a single query, allowing the model to see the scores of other elements.
-The approach corresponds to pointwise ranking approach Sun et al.
-(2023); Sachan et al.
-(2022) in the RAG literature.
-VA approaches correspond to pairwise ranking approaches Qin et al.
-(2024) in the RAG literature, but our contribution focuses on leveraging the theoretical properties of classical algorithms, which has been undermined in the literature.
-VA Powersort overrides Python’s __gt__ operator with an LLM call and applies the built-in sorted function, which uses Powersort (Munro and Wild, 2018) that runs $O(n\log n)$ comparisons worst-case, the best runtime complexity achievable by comparison sort.
-VA Bitonic sorting network uses the Bitonic sorting network (Batcher, 1968) from parallel sorting literature to improve the throughput.
-It performs $O(n(\log n)^{2})$ comparisons, but runs in $O((\log n)^{2})$ time due to parallelism.
-As an additional optimization, we use the optimal sorting network (Knuth, 1997; Bundala and Závodnỳ, 2014) when the input size is 16 during its recursion.
-Robust VAs (Powersort/Bitonic) make individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting.
-For an odd $K$, if the oracle is correct with probability $p$, the voting result is correct with at least probability $1-e^{-K(2p-1)^{2}/2}$ (Emamjomeh-Zadeh and Kempe, 2018) due to Hoeffding bound (Hoeffding, 1963).
-Finally, Probabilistic VAs (Powersort/Bitonic) instead rely on comparing the probabilities / logits over the yes/no answers to perform each comparison query.
-While the naive probabilistic VA is equivalent to the naive VA with greedy sampling, we additionally have Symmetrized VA to address two biases in LLM: Positional bias and psycophancy (the tendency to reply “yes”) .
-Let $p_{12}$ and $p_{21}$ be next-token probabilities for answering “yes” with the original prompt and a prompt whose $x_{1}$ and $x_{2}$ are swapped.
-Due to the bias, $p_{12}\not=1-p_{21}$.
-Symmetrized VA tests $\frac{p_{12}+(1-p_{21})}{2}>\frac{(1-p_{12})+p_{21}}{2}$ as a comparison oracle.
 Evaluation
@@
 The strength of VA is most pronounced with the small 1.7B model, where other approaches failed.
-Notably, VAs with a 1.7B model outperformed the baseline with a 32B model, clearly showing the advantage of using the sorting algorithms developed in the literature.
+Notably, VAs with a 1.7B model outperformed the baseline with a 32B model, clearly showing the advantage of using the sorting algorithms developed in the classical discrete algorithm literature.
 Finally, Table 4 compares the runtime of each approach.
@@
 This shows that VAs can directly benefit from the theoretical results in the parallel sorting literature.
-Note that the parallel speedup in bitonic sort does not reach the theoretical limit (x11.4; 171 comparators and 15 steps for 25 inputs) because the maximum batch size is limited by the GPU; multi-GPU experiments with data-parallelism would improve this.
-The Effect of Prefix Caching
-Each query in VA sort requires a single next-token prediction.
+Note that the parallel speedup in bitonic sort does not reach the theoretical limit (x11.4; 171 comparators and 15 steps for 25 inputs) because the maximum batch size is limited by the GPU; multi-GPU experiments with data-parallelism would improve this result even further.
+Effect of Prefix Caching
+Each query in VA sort requires a single next-token prediction for yes/no.
 In Transformer architecture, it requires $O(t^{2})$ computations for $t$ input tokens.
-However, State-of-the-Art inference engines such as vllm Kwon et al.
-(2023) and SGLang Zheng et al.
-(2024) cache the prefix kv-cache over multiple queries.
+However, State-of-the-Art inference engines such as vllm (Kwon et al., 2023) and SGLang (Zheng et al., 2024) cache the prefix kv-cache over multiple queries.
 This is particularly helpful in reducing the inference cost of VA-style queries majority of which share the prefix.
@@
 In the exhaustive pairwise comparison among $n$ strings ($O(n^{2})$ comparisons, equivalent to insertion sort etc.), a naive inference engine requires $O((L_{p}+2L)^{2})$ runtime per each yes/no query, thus $O(n^{2}(L_{p}+2L)^{2})$ runtime in total, without parallelism.
-Assume that an input consists of $N+M$ tokens (prefix and suffix), and the kv-cache of $N$ prefix tokens is already available.
+We analyze the runtime reduction due to kv-caches.
+Assume that an input query consists of $N+M$ tokens (prefix and suffix), and the kv-cache of $N$ prefix tokens is already available.
 Each $N+t$-th new input token attends to N+(t1)N+(t-1) existing tokens, which makes $N+M+1$-th token require $\sum_{t=1}^{M}(N+(t-1))=O(NM+M^{2})$ runtime.
-With this in mind, assuming that the fixed prompt already has a kv-cache, the runtime of each comparison query reduces to $O(2L_{p}L+4L^{2})$, totalling $O(L_{p}^{2}+n^{2}(2L_{p}L+4L^{2}))$ runtime.
+In other words, the runtime of each comparison query reduces to $O(2L_{p}L+4L^{2})$, totalling $O(L_{p}^{2}+n^{2}(2L_{p}L+4L^{2}))$ runtime.
 Moreover, since $x_{1}$ is repeated $n$ times over $n^{2}$ comparisons, it further reduces to $O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}((L_{p}+L)L+L^{2}))=O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}(L_{p}L+2L^{2}))$.
-This is significant because the prompt tends to be much longer than each string.
-Table 4 demonstrates the empirical speedup due to kv-cache in two State-of-the-Art inference engines. RadixAttention in SGLang is particularly helpful for VAs whose queries share large prefixes.
+Although we assumed a simplified, exhaustive ($O(n^{2})$) comparison scenario, this speedup is significant in practice.
+Table 5 demonstrates the empirical speedup due to kv-cache in two state-of-the-art inference engines.
+RadixAttention in SGLang is particularly helpful for VAs whose queries share large prefixes.
 5 Verbalized Clustering
-Verbalized Clustering attempts to cluster a set of strings $S$ into a family of subsets $\mathcal{S}$ based on their similarity.
-The problem with text-based clustering is that the distance function for natural language strings is difficult to define.
-All approaches to this task are embedding methods.
-Method, dataset, and prompt details are available in the appendix Sec.
-B.
-All results are averages of 5 random seeds.
-Jaccard Baseline creates a distance matrix using Jaccard distance of token sets of strings, then applies sklearn’s MDS (Kruskal, 1964) to create an embedding.
-Sentence Transformer (ST) Baseline uses Qwen3-8B-Embedding directly, which was selected as the top performer in the MTEB leaderboard (Muennighoff et al., 2022).
-VA t-STE: This is a VA based on Student-t Stochastic Triplet Embedding (Van Der Maaten and Weinberger, 2012, t-STE), which makes triplet comparison queries to LLM.
-Triplet comparison takes three elements $x$, $y$ and $z$ in a metric space $\|\cdot\|$ and query if $\|x-y\|<\|x-z\|$.
-We ask the LLM a question such as “Is X closer to Y than is to Z?
+Topic clustering attempts to cluster a set of strings $S$ into a family of subsets $\mathcal{S}$ based on their similarity.
+The challenge with text-based clustering is defining a metric space and a distance function.
+While traditional approach toward this task performs $k$-Means or other standard clustering algorithms on the distance matrix over the vector embeddings of the input strings, Verbalized Clustering instead leverage triplet comparison queries to LLM.
+Triplet comparison takes three elements $x$, $y$ and $z$ in a metric space $\|\cdot\|$ and query if $\|x-z\|<\|y-z\|$, i.e., we ask the LLM a question such as “Is Z similar to X than is to Y?
 X: $x$, Y: $y$, Z: zz”.
-For each $x\in S$, we randomly sample $k$ queries from $S^{3}$ and their results from an LLM, then generate $D$-dim embeddings by minimizing the number of violations to the results.
-We show the results with hyperparameters $k\in{\left\{100,200\right\}},D=8$.
-We evaluated the methods across four datasets.
+We propose Verbalized Clustering (Alg.
+1), which combines $O(n\log n)$ approximate Delaunay graph construction (Haghiri et al., 2017), a triplet-based algorithm which was originally developed for $k$-Approximate Nearest Neighbor Search, and $O(n\log^{2}n)$ Clauset-Newman-Moore greedy modularity maximization (Clauset et al., 2004), an algorithm that clusters the set of nodes on a graph based on how densely connected the nodes are.
+The latter does not involve LLM calls, thus $O(n\log n)$ with a large constant factor for LLM calls dominates the runtime rather than $O(n\log^{2}n)$.
+1:function Cluster(strings $S$, int $k$, int $n$):
+2: Initialize a graph $G=(V,E)\leftarrow(S,\emptyset)$
+3:for $0<i<k$ do
+4:  Roots $x,y\leftarrow$ randomly chosen from $S$
+5:  $E\leftarrow E\cup{\left\{(x,y)\right\}}$, $S^{\prime}\leftarrow S\setminus{\left\{x,y\right\}}$
+6:  if $S^{\prime}\not$$\emptyset$ then
+7:   $\textsc{Recurse}(x,{\left\{z\in S^{\prime}\mid\|x-z\|<\|y-z\|\right\}})$
+8:   $\textsc{Recurse}(y,{\left\{z\in S^{\prime}\mid\|x-z\|>\|y-z\|\right\}})$
+9:return greedy_modularity_communities($G,n$)
+10:function Recurse(parent $p$, strings $S$):
+[…]
+[…]
+13:if $S^{\prime}\not$$\emptyset$ then
+14:  $\textsc{Recurse}(x,{\left\{z\in S^{\prime}\mid\|x-z\|<\|y-z\|\right\}})$
+15:  $\textsc{Recurse}(y,{\left\{z\in S^{\prime}\mid\|x-z\|>\|y-z\|\right\}})$
+Algorithm 1 Verbalized Clustering with strings $S$, int $k$, number of output clusters $n$
+Table 6: Adjusted Rand Index across methods and benchmarks. The top results are highlighted in bold.
+$k$ Medarxiv xkcdcolors-h xkcdcolors-l Qwen3-8B 20 0.16 0.42 0.10 50 0.14 0.39 0.06 Qwen3-14B 20 0.14 0.51 0.10 50 0.18 0.42 0.12 Qwen3-32B 20 0.17 0.46 0.11 50 0.20 0.46 0.16 granite-embedding-125m-english 0.13 0.18 0.00 Qwen3-Embedding-8B 0.11 0.43 0.01 all-MiniLM-L6-v2 0.11 0.19 0.01
+We compare VA clustering with Sentence Transformer (ST) Baseline, which uses Qwen3-8B-Embedding directly, which was selected as the top performer in the MTEB leaderboard (Muennighoff et al., 2022).
+The embedding vectors are then processed by K-Means to obtain the cluster labels.
+We evaluated the methods across three datasets.
 Medarxiv (Muennighoff et al., 2022) contains the titles and the categories of medical papers on MedArxiv.
 The category forms a cluster.
-IT Field Defects consists of $78$ technical IT field defect reports submitted to our internal defect tracking system (unpublished), where some reports are manually marked as duplicates of another, forming a cluster.
-xkcdcolors-h/l are two datasets of 100 colors sampled from (Munroe, 2010).
+xkcdcolors-h/l are two datasets of 100 colors we created from (Munroe, 2010).
 Each entry has a non-standard color name (e.g., baby blue, vomit) and a corresponding RGB value.
-The difference between h and l variants is that we defined 6 ground truth clusters with sklearn.AgglomerativeClustering using a lexicographic distance that prioritizes hue and lightness in the HSL color model, respectively.
+The difference between h and l variants is that we defined 6 ground truth clusters with sklearn.AgglomerativeClustering using a lexicographic distance that prioritizes hue and lightness in the ground-truth HSL color values, respectively.
 Assuming that colors tend to be more strongly associated with hues, xkcdcolors-l represents an out-of-distribution (OOD) task.
-We evaluate the embeddings with a score defined in Eq.
-1, which is the average rank of the members of the cluster each string belongs to, when sorted by the Euclidean distance.
-The score is normalized to $[0,1]$, where 1 indicates that cluster members are embedded closer.
-Given the input $S$ and a string $s\in S$, let $C(s)$ be the ground-truth cluster such that $s\in C(s)$.
-Let $i=\textsc{rank}(s,s^{\prime})\in{\left\{0,1,\ldots|S|-1\right\}}$ be an integer such that, when sorted by the pairwise Euclidean distance $d(e(s),e(s^{\prime\prime}))$ between $s$ and a string $s^{\prime\prime}$ across $s^{\prime\prime}\in S$ using embedding $e$, then $d(e(s),e(s^{\prime}))$ ranks $i$-th.
-The score is defined as:
-score $=\frac{1}{|S|}\sum_{s\in S}\frac{1}{|C(s)|-1}\sum_{s^{\prime}\in C(s)\setminus{\left\{s\right\}}}1-\frac{\textsc{rank}(s,s^{\prime})}{|S|}\in[0,1]$ (1)
+We evaluate the accuracy using the Adjusted Rand Index (Hubert & Arabie, 1985).
 Table 6 shows the results.
-Jaccard fared surprisingly well in IT Defects, where the input often contains long output logs whose tokens carry significant information, but it underperformed in tasks that require semantic comprehension.
-Sentence Transformer achieves strong performance in the standard clustering tasks such as Medarxiv and IT Defects.
+Sentence Transformer achieves strong performance in the standard clustering tasks such as Medarxiv.
 However, it fell short in xkcdcolors-h/l that requires information about the distance criteria.
-The score 0.53 in the xkcdcolors-l is close to 0.5, i.e., its distance ranking is only marginally better than a dice roll.
-VA t-STE outperformed traditional baselines with $k=200$.
+The scores in the xkcdcolors-l are close to 0, i.e., its output clusters is almost equivalent to a dice roll.
+In all datasets, Verbalized Clustering outperformed traditional baselines.
 We highlight its strong generalization in OOD tasks where other approaches failed, due to its strong algorithmic bias combined with flexible LLM queries.
+6 Verbalized Submodular Maximization
+Retrieval Augmented Generation (RAG) is a mainstream approach for reducing LLM hallucinations by retrieving documents that are relevant to the query from a large database and appending such documents to the prompt so that LLMs can draw answers from the documents.
+While existing work tend to simply rank the documents via vector embedding similarity search, this approach fails when the task requires combining information in multiple documents to draw conclusions, because all the top-ranked documents may be similar to each other and lack crucial information.
+To address this limitation, existing work formulated multi-hop Q&A tasks (Yang et al., 2018) and observed that there are several types in each ground-truth retrieval (e.g., bridging, comparison, etc), differing in how they chain multiple information together.
+Existing work proposed various heuristic methods that additionally require fine-tuning (Park et al., 2025), based on the observations.
+On the other hand, traditional embedding-based approaches to multi-hop reasoning rely on a classical notion of submodularity: Given a search space $C$ and a solution $S\subseteq C$, they assume the evaluation function $f(S)$ has diminishing returns, i.e., $\forall x\in C,S\subseteq S^{\prime}\subseteq C;f(S^{\prime}\cup{\left\{x\right\}})-f(S^{\prime})\leq f(S\cup{\left\{x\right\}})-f(S)$.
+This approach, however, only works in a setting that has a well-defined evaluation function $f$ that has submodularity.
+We propose Verbalized Greedy Submodular Maximization (VGSM), a verbalized algorithm that follows the classical greedy algorithm that guarantees 11/e1-1/e optimality.
+While the notion of submodularity requires a concrete evaluation function $f$, VGSM specifies a implicit score function by giving a prompt shown in Fig.
+1 to the LLM.
+Although LLMs are not able to produce the absolute score value $f(S\cup{\left\{x\right\}})$ when it selects a new element $x\in D\setminus S$, this is not an issue in greedy submodular maximization because it only requires a relative, contrastive measure that encodes the selection policy / preference, which originates from LLMs’ Plackett-Luce reward model.
+[…]
+Figure 1: The submodularity-inducing query for VGSM. Note that $X$ contains all non-selected documents $D\setminus S$.
+We evaluated the approach using HotpotQA (Yang et al., 2018) (fullwiki, hard) and CUAD (Hendrycks et al., 2021) datasets.
+Each dataset contains a set of 3-tuples $(Q,A,D)$ where $D$ is a set of documents, $Q$ is the question and $A$ is the ground-truth answer.
+We assume a generative model $p(A|D,Q)=\sum_{S}p_{\phi}(A|S,Q)p_{\psi}(S|Q,D)$, where $S$ is the documents selected by the retrieval model $\psi$, and $p_{\phi}(A|S,Q)$ represents the likelihood of the answer by an answering model $\phi$.
+Given $\psi$ and $\phi$, we evaluate the performance of $\psi$ by $\log p_{\phi}(A|S,Q)=\sum_{t}\log p_{\phi}(A_{t}|A_{1:t-1},S,Q)$, i.e., the sum of log likelihoods over each token in the answer string $A$.
+This log likelihood is obtained by feeding the LLM a prompt concatenating $Q,S,A$ in order.
+We compared VGSM with seven baselines.
+TF-IDF and BM25 are the traditional bag-of-words baselines that provide a similarity metric $f(x,y)$.
+It sorts documents $D$ by $f(Q,d)$ for $d\in D$, and selects $k$ most similar documents.
+Embedding baseline uses state-of-the-art BERT/Transformer embedding models from SentenceTransformers library.
+Let $e(x)$ be the embedding vector of a text string $x$.
+It sorts $D$ by the cosine distance between $e(Q)$ and $e(x)$ for $x\in D$, and selects the top-$k$ instances (closest to $Q$).
+MMR-$f$ baselines applies Maximum Marginal Relevance (MMR) rule (Carbonell & Goldstein, 1998) to the similarity metric $f$.
+(When $f$ is a distance, use the additive inverse.) It converts $f$ into a non-monotone submodular function $f_{MMR}(x,S)=\lambda f(x,Q)-(1-\lambda)\max_{y\in S}f(x,y)$, where $S$ is a set of already selected documents, then apply the greedy submodular maximization algorithm.
+We evaluate this baseline to assess the impact of submodularity on non-LLM-based scoring functions.
+However, since MMR is non-monotone, it loses the 11/e1-1/e optimality guarantee.
+Finally, Verbalized Sorting uses the Symmetrized Powersort discussed in Sec.
+4 to sort $D$ and retrieve top-$k$ elements.
+We evaluate this in order to see the impact of using LLMs for multi-hop retrieval task without imposing submodularity.
+Using each method, we retrieved $k=3,6,12$ documents and evaluated $\textsc{logp}=\log p_{\phi}(A|S,Q)$.
+We varied $\phi$ and $\psi$ across Qwen3 models of various sizes ($\phi\in{\left\{\text{0.6B},\text{32B}\right\}}$, $\psi\in{\left\{\text{0.6B},\text{4B},\text{32B}\right\}}$).
+For a faster evaluation, VGSM and Verbalized Sorting pre-selects $2k$ or $4k$ documents using the embedding baseline before running the respective algorithm.
+Table 7 shows logp of various methods.
+Traditional TF-IDF and BM25 baselines consistently performed worse than the embedding model, as expected.
+MMR did not help TF-IDF, BM25 or the embedding models, which could be attributed to the lack of monotone submodularity in the MMR scheme.
+For every embedding model $e$, combination with VGSM ($|D|\rightarrow{\left\{2k,4k\right\}}\rightarrow k$ documents) consistently achieved higher logp than simply selecting $k$ documents with $e$, demonstrating the effectiveness of VGSM.
+Verbalized Sorting performed worse than VGSM, which demonstrates the impact of the implicit submodular objective encoded by the LLMs.
+On the performance of VGSM, we observed: (1) Regardless of the method, allowing a larger $k$ tends to improve logp because it increases the chance of successful retrieval.
+(2) Regardless of the method, using a larger answer model $\phi$ improved logp, but it did not affect the relative performance among the retrievers $\psi$.
+(3) When $k=12$ and pre-selecting $4k=48$ documents, VSGM with $\psi=$ Qwen3-0.6B performed worse than with $2k=24$ or with smaller $k$ (highlighted in red).
+This is likely because selecting from 48 candidates challenged the long context comprehension of the model.
+A possible mitigation to this issue is replacing the $4k/2k$ selection step with Verbalized Maximum (Sec.
+3), whose context only considers 2 documents at a time.
+Table 7: Comparing logp over various retrieval models.
+In all cases, the embedding model that (pre-)selected kk/2k2k/$4k$ elements is Qwen3-embedding-8B.
+All other models are implied to be the number of parameters of Qwen3-family of models.
+We did not perform an exhaustive experiments over all parameters which leaves some cells empty, however our main claim holds.
+hotpotQA CUAD $\phi=$ 0.6B $\phi=$ 32B $\phi=$ 0.6B $\phi=$ 32B $\psi$ $k=3$ 6 12 3 6 12 3 6 12 3 6 12 tfidf -10.6 -9.9 -9.1 -5.9 -5.4 -5.0 -309.4 -289.5 -269.0 -206.6 -191.3 -177.0 tfidf+MMR $\lambda=0.2$ -11.0 -11.0 -10.8 -6.2 -6.2 -6.2 -326.3 -325.4 -320.9 -219.3 -217.6 -211.6 tfidf+MMR $\lambda=0.5$ -10.6 -9.9 -9.4 -6.0 -5.7 -5.4 -319.2 -308.1 -299.9 -212.6 -202.8 -196.5 tfidf+MMR $\lambda=0.8$ -10.4 -9.7 -9.1 -5.8 -5.3 -5.0 -310.6 -289.6 -268.7 -206.9 -192.0 -177.1 bm25 -9.9 -9.7 -9.1 -5.4 -5.2 -5.0 -308.1 -288.3 -267.6 -205.6 -190.5 -175.7 bm25+MMR $\lambda=0.2$ -10.7 -10.6 -10.3 -6.0 -5.9 -5.7 -317.6 -315.6 -311.8 -212.0 -208.7 -204.4 bm25+MMR $\lambda=0.5$ -10.4 -10.1 -9.8 -5.9 -5.7 -5.5 -314.1 -307.1 -301.6 -208.7 -202.6 -197.4 bm25+MMR $\lambda=0.8$ -10.0 -9.6 -8.8 -5.6 -5.3 -4.9 -309.1 -295.3 -280.9 -205.9 -195.1 -185.1 embed(top-$k$) -9.6 -9.1 -8.6 -5.2 -5.1 -5.1 -291.0 -269.2 -253.9 -197.6 -180.9 -168.7 embed(top-kk)+MMR $\lambda=0.2$ -10.7 -10.6 -10.3 -5.7 -5.8 -5.7 -310.4 -305.9 -296.1 -208.9 -202.7 -193.8 embed(top-kk)+MMR $\lambda=0.5$ -9.7 -9.4 -8.6 -5.4 -5.4 -5.1 -305.8 -292.9 -279.1 -206.1 -194.6 -183.2 embed(top-kk)+MMR $\lambda=0.8$ -9.5 -9.2 -8.4 -5.2 -5.1 -5.0 -294.3 -276.1 -255.3 -199.2 -185.1 -169.2 embed(top-2k2k)+VA sort 0.6B -10.3 -9.9 -10.0 -301.5 -284.1 embed(top-4k4k)+VA sort 0.6B -11.0 -11.3 -10.7 -307.4 embed(top-2k2k)+VGSM 0.6B -9.6 -9.1 -8.6 -5.2 -5.1 -5.0 -290.5 -268.6 -251.3 -196.4 -180.4 -167.1 embed(top-4k4k)+VGSM 0.6B -9.5 -9.2 -10.0 -5.3 -5.2 -5.7 -289.8 -267.6 -250.3 -196.8 -179.6 -165.7 embed(top-2k2k)+VGSM 4B -9.2 -8.7 -8.2 -280.3 -267.3 -248.6 embed(top-4k4k)+VGSM 4B -9.0 -8.6 -8.2 -282.6 -265.0 -241.5 embed(top-2k2k)+VGSM 32B -9.2 -8.5 -7.9 -277.9 -261.7 -240.3 embed(top-4k4k)+VGSM 32B -8.6 -8.1 -7.8 -273.9 -250.4 -228.9
 7 Related Work
@@
 Formalization approach to planning is also called a parse-and-solve model (Collins et al., 2022) as opposed to a LLM-as-Planner model that naively generates a plan with an LLM.
-Thought-of-Search (Katz et al., 2024) simulates a systematic search algorithm, which could be seen as an instance of VAs.
-To our knowledge, there is no work that simulates traditional graph search algorithms (e.g., A* (Hart et al., 1968), GBFS (Bonet and Geffner, 2001)) in a VA manner, i.e., by replacing all of its atomic operations (e.g., successor generation, heap-based priority queues, hash computation for duplicate detection) with LLM queries, although some work replace some atomic operations, especially successor generations (Katz et al., 2024).
-A number of recent papers in the Algorithm with Predictions (Mitzenmacher and Vassilvitskii, 2022, ALPS) community achieved a significant success in augmenting algorithms by data-driven oracles.
+To our knowledge, there is no work that simulates traditional graph search algorithms (e.g., A* (Hart et al., 1968), GBFS (Bonet & Geffner, 2001)) in a VA manner, i.e., by replacing all of its atomic operations, e.g., successor generation, heap-based priority queues, hash computation for duplicate detection, with LLM queries.
+Some work replaces some atomic operations through code generation: successor function and goal condition in Thought of Search (Katz et al., 2024) and collision checking in STPR (Djuhera et al., 2025).
+A number of recent papers in the Algorithm with Predictions (Mitzenmacher & Vassilvitskii, 2022, ALPS) community achieved a significant success in augmenting algorithms by data-driven oracles.
 The particular strength of this line of work is the rigorous theoretical analysis of the error rate.
 VAs is a form of ALPS specialized in LLM-based oracles, thus combining ideas from ALPS community with VAs is a promising direction for future work.
-Combining data-driven heuristics with existing algorithms in the classical AI tasks is also a common practice in the recent literature (e.g., classical planning Núñez-Molina et al.
-(2024), boolean SAT Duan et al.
-(2020), logistics tasks Xin et al.
-(2021)).
-6 Conclusion, Discussion and Future Work
+Combining data-driven heuristics with existing algorithms in the classical AI tasks is also a common practice in the recent literature (e.g., classical planning (Núñez-Molina et al., 2024), boolean SAT (Duan et al., 2020), logistics tasks (Xin et al., 2021)).
+8 Conclusion and Future Work
 We presented Verbalized Algorithms (VAs), a disciplined paradigm for performing algorithmic reasoning with LLMs.
@@
 It also leverages their established theoretical results to improve the runtime of such reasoning.
-While this paper explored only a limited number of applications due to space, we are already pursuing several exciting verbalizations, including submodular maximization and pareto-front computation for diverse / non-dominated document retrieval, DPLL for combinatorial reasoning, and median/top-$k$ selection, etc.
-The potential of this paradigm is large, and we invite various ideas for unique verbalizations and new applications.
-For example, can Floyd’s tortoise-and-hare cycle detection algorithm for linked lists be verbalized to perform a tautology detection?
-Garbage collection algorithms to detect unimportant sentences in a paper draft?
+While this paper explored only a limited number of applications due to space, we are already pursuing several exciting verbalizations, including pareto-front computation for diverse / non-dominated document retrieval, DPLL for combinatorial reasoning, and median/top-$k$ selection, etc.
+Can Floyd’s tortoise-and-hare cycle detection algorithm for linked lists be verbalized to perform a tautology detection?
+Can garbage collection algorithms detect unimportant sentences in a paper draft?
+The potential of this paradigm is large and we invite various ideas for unique verbalizations and new applications.
 Appendix
@@
 A.2 Kendall-Tau Score
-The Kendall Tau correlation coefficient Kendall (1945) is a non-parameteric statistic that measures how much order is preserved between two sets of rankings.
+The Kendall Tau correlation coefficient (Kendall, 1945) is a non-parameteric statistic that measures how much order is preserved between two sets of rankings.
 Kendall-Tau score has several ways to compute ($\tau_{a},\tau_{b},\tau_{c}$) depending on the existence of ties or the nature of the observations.
@@
 where $|T_{i}\cap T_{j}|$ represents the cardinality of the intersection and $|T_{i}\cup T_{j}|$ represents the cardinality of the union.
-B.2 t-STE
-Van Der Maaten and Weinberger (2012) proposed a heavy-tailed kernel to measure local similarities between data points.
-In particular, we opt to use a Student-$t$ kernel with $\alpha$ degrees of freedom by defining:
-$p_{ij\ell}=\frac{\left(1+\frac{\lVert x_{i}-x_{j}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}}{\left(1+\frac{\lVert x_{i}-x_{j}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}+\left(1+\frac{\lVert x_{i}-x_{\ell}\rVert^{2}}{\alpha}\right)^{-\tfrac{\alpha+1}{2}}}$
-We refer to the resulting technique as t-Distributed STE (t-STE). Our implementation sets $\alpha=D-1$, where $D$ is the number of dimensions in the resulting embedding.
-The formulation of the triplet probabilities is motivated by the success of unsupervised dimensionality reduction technique that also employ heavy-tailed similarity kernels (Maaten and Hinton, 2008).
-B.3 LLM Query Prompts
+B.2 LLM Query Prompts
 Triplet comparison query uses the following shared prompt, where criteria is a dataset-specific prompt.
@@
 criteria = f"You will be given three paper titles X, Y, and Z. Does the paper Z discuss a topic that is closer to the topic in X than to the topic in Y?"
-IT Defect dataset:
-criteria = f"You will be given three defect reports X, Y, and Z, sent by the customers of our GPFS filesystem storage system.
-They may contain stack traces of various programs (python, kernel drivers, ...).
-Similar repetitive lines are omitted as [...].
-Considering the cause of the defects (i.e.
-ignoring timestamp/IP/MAC addresses), are X and Z more similar than Y and Z are?
-"
 xkcdcolors-h dataset:
@@
 "
-B.4 IT Defect Reports Dataset
-Let $D$ denote the set of all defects.
-A subset $D^{\prime}\subseteq D$ was manually labeled as duplicates of several other defects in $D$.
-We selected 78 data points so that the final dataset contains 25 such clusters, each of size at least 2.
-Example Datapoint (redacted/simplified)
-The following is an example of a redacted/simplified defect record that appear in the dataset.
-  "id": 347852,
-  "description":
-  "System login:
-    EMS - X.XX.XXX.XXX root/cluster
-    S6K w/ Falcon attached - lothal-qa6-[1 && 2] Code levels:
-    S6k - 6230_B19
-    FCM4 - 4_3_3
-    Falcon - 5342 --> 5343
-   Snap location:
-    glogin:/u/DUMPS/R.347852/
-   What I was doing:
-    I was upgrading Falcon enclosure from 5342 to 5343.
-   What I expected:
-    ’XXXXfirmware’ will start and the Falcon will go through the upgrade process.
-   What happened:
-    ’mmchrimware’ completed but there are errors printed.
-    The Falcon does appear to have the updated firmware.
-    However, there are issues when trying to get a print
-    from the Left module when running ’essenclosures’.
-   Analysis:
-    sg91:13:/dev/sg91::naa.500093d006585000/10910::[1.0.91.0]:/dev/sg91:XXX:5149-091:5342
-    [10:48:31.258608008] XXXXXXFirmwareVer2 checking if enclosure needs to be updated
-    [10:48:31.260567945] XXXXXXFirmwareVer2 firmware file found
-    [10:48:31.262950549] XXXXXXFirmwareVer2 -> ...
-    [10:48:31.350737778] XXXXXXFirmwareVer2 ...
-    [10:48:31.362309567] XXXXXXFirmwareVer2 lothal-qa6-1-hs: [E] Error 1 updating firmware ...
-    sg_ses failed: Transport error
-  "duplicate_of": [],